# DeepVoid: A Deep Learning Void Detector

### v1

@@ Abstract
We present DeepVoid, an application of deep learning trained on a physical definition of cosmic voids to detect voids in density fields and galaxy distributions. By semantically segmenting the IllustrisTNG simulation volume using the tidal tensor, we train a deep convolutional neural network to classify local structure using a U-Net architecture for training and prediction. The model achieves a void F1 score of 0.96 and a Matthews correlation coefficient over all structural classes of 0.81 for dark matter particles in IllustrisTNG with interparticle spacing of $\lambda=0.33$ $h^{-1}$ Mpc. We then apply the machine learning technique of curricular learning to enable the model to classify structure in data with significantly larger intertracer separation. At the highest tracer separation tested, $\lambda=10$ $h^{-1}$ Mpc, the model achieves a void F1 score of 0.89 and a Matthews correlation coefficient of 0.6 on IllustrisTNG subhalos.
@@
revtex4-1Repair the float
1 Introduction
The large-scale structure of the universe, as observed in galaxy redshift surveys, resembles a web-like structure (Bond et al., 1996). This structure is, by volume, chiefly comprised of cosmic voids, underdense regions that span tens of megaparsecs and are bounded by sheets and filaments of galaxies that flow into large galaxy clusters. Depending on the definition, voids fill approximately $\sim$70of the volume of the universe (Pan et al., 2012). Voids were found in some of the earliest galaxy and cluster redshift surveys (e.g., Gregory & Thompson, 1978; Jõeveer et al., 1978) and were shown to be ubiquitous features of the large-scale structure by dense three-dimensional galaxy redshift surveys (de Lapparent et al., 1986; Geller & Huchra, 1989; Shectman et al., 1996).
Voids represent an interesting opportunity for cosmology and physics (Cai & Neyrinck, 2025). Voids grow from initial troughs in the primordial field of density fluctuations and continue to become more vacuous as the large-scale structure in the Universe evolves. They are ideal laboratories for studying modifications to General Relativity (Perico et al., 2019) by observing, for example, the weak lensing signature of voids (Baker et al., 2018) or the gravitational redshift of Cosmic Microwave Background (CMB) photons by large voids, known as the Integrated Sachs-Wolfe (ISW) effect (Cai et al., 2014).
Investigations of dark energy (DE) can also greatly benefit from studying voids. Because overdense structures such as galaxy clusters formed long before dark energy could strongly influence the evolution of large-scale structure, voids are the primary component of the cosmic web influenced by dark energy. Voids act as super-Hubble bubbles that expand faster than the Hubble flow (Goldberg & Vogeley, 2004). To leading order, the accelerated expansion of voids is the primary signature of dark energy; the overdense structures accelerate away from each other because the voids are expanding between them (Sheth & van de Weygaert, 2004).
Several specific cosmological tests using cosmic voids are of particular interest and motivate how our void finder is designed and implemented. The shapes of voids provide an avenue to test cosmological parameters through the Alcock-Paczyński (AP) effect (e.g., Alcock & Paczynski, 1979; Lavaux & Wandelt, 2012; Hamaus et al., 2020). The size spectrum of voids is sensitive to both the growth of structure and the dark energy equation of state (e.g., Nadathur, 2016; Verza et al., 2019; Contarini et al., 2022a, 2023). Observation of the ISW effect from voids can further constrain cosmological models (e.g., Granett et al., 2008; Kovács, 2018).
One of the chief difficulties in finding and using voids for cosmology and astrophysics is the definition of a void. There is no universally agreed-upon definition, but studies generally rely on finding underdensities in the galaxy distribution, with methods mostly varying in the shape of the density estimation kernel. Popular void-finding methods for galaxy surveys utilize techniques that include the detection of voids as spherical regions or unions of multiple spheres (e.g., Kauffmann & Fairall, 1991; El-Ad & Piran, 1997; Hoyle & Vogeley, 2002; Banerjee & Dalal, 2016; Zhao et al., 2016; Paz et al., 2023), and watershed segmentation of the galaxy density field (e.g., Platen et al., 2007; Neyrinck, 2008; Sutter et al., 2015; Nadathur et al., 2019). Void catalogs constructed from identical galaxy samples can differ drastically based on how a void is defined in that study (Colberg et al., 2008). These different definitions make it complicated to compare analyses based on different void catalogs. Many of these methods also rely to varying degrees on fine-tuned parameters and different a priori notions of what constitutes a void and how they should appear.
In this work, we demonstrate a proof of concept of an alternative approach: define voids in terms of the physics that creates and shapes them, and train a deep neural network to recognize the spatial signature of that physical process. Here we describe a physics-based void detector based on convolutional neural networks (CNNs) called DeepVoid. By training CNNs on physical definitions of a void, we can obtain more targeted void catalogs for different cosmological tests of interest. Previous investigations of using deep learning for morphological classification include Aragon-Calvo (2019) and Inoue et al. (2022). Unlike previous work, DeepVoid is targeted to the specific task of detecting voids in galaxy redshift surveys, a task which grows more challenging as these surveys probe deeper into the universe.
The eventual goal of this approach is to fully leverage the cosmological information available in current and upcoming galaxy surveys and instruments, such as the Dark Energy Spectroscopic Instrument (DESI; DESI Collaboration et al., 2016), Vera C. Rubin Legacy Survey of Space and Time (LSST; LSST Science Collaboration et al., 2009), Euclid (Euclid Collaboration et al., 2024), Nancy Grace Roman Space Telescope (Schlieder et al., 2024), and Subaru/Hyper Suprime Camera (HSC; Miyazaki et al., 2018). These surveys and others will provide maps of large-scale structure to unprecedented depths and with remarkable precision. They will probe ranges of cosmological redshift over which the evolution of voids is apparent and over which details of DE properties, possible modifications to General Relativity, and neutrino properties might be probed (Massara et al., 2015; Pisani et al., 2015; Perico et al., 2019; Avsajanishvili et al., 2024).
Important challenges include that the galaxy catalogs from these new surveys may be quite sparse compared to more local galaxy surveys and that large-scale structure is probed by many different types of galaxy tracers with different biases (Euclid Collaboration et al., 2022; Wang et al., 2022; Hahn et al., 2023a; Zhou et al., 2023; Raichoor et al., 2023; Chaussidon et al., 2023). The effect of “tracer bias” on voids has been shown to affect the void size distribution (Nadathur & Hotchkiss, 2015; Contarini et al., 2019), impact the tracer profile vs. matter profile in voids (Pollina et al., 2017), and alter the velocity profile in voids, which is important for using void-galaxy correlations to study redshift space distortions (RSD; Massara et al., 2022). Current methods of estimating the void spectrum involve corrections for that bias applied after making a void catalog, based on analysis of mock galaxy catalogs (Contarini et al., 2022b). We will minimize the need for such post-detection void corrections by using mock catalogs to train variants of DeepVoid for specific galaxy target classes. In this proof-of-concept paper, we demonstrate an approach to tackling the sparse sampling problem. In a follow-up paper, we will examine how the deep learning approach addresses the variety of galaxy tracers.
In this work, we present DeepVoid, a novel multi-class method of segmenting the morphology of structure in cosmological simulations using CNNs trained to match the classes assigned from properties of the gravitational potential. This paper is organized as follows. In Section 2, we describe the simulations to which our physical definition of a void is applied to define our truth table. The physical definition itself is discussed in Section 3. Aspects of deep learning and the U-Net architecture are detailed in Section 4. Model training and performance are presented in Sections 5 and 6, respectively. Training and performance of models that predict on sparse samples of halos are described in Section 7. Conclusions and thoughts about future work are shared in Section 8. Additional details on multi-class performance metrics, our use of batch normalization, and different loss functions can be found in the Appendices.
2 Simulation Data
2.1 Illustris TNG
The Next Generation Illustris Simulation (TNG) is a suite of cosmological simulations released in 2018 (Springel et al., 2018; Pillepich et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018) and is the successor to the original Illustris simulation (Vogelsberger et al., 2014). TNG is a gravo-magnetohydrodynamical simulation run using the moving-mesh code AREPO (Springel, 2010). TNG is made up of three volumes of different sizes: 35, 75, and 205 $h^{-1}$ Mpc on a side. All TNG runs were computed with the Planck 2015 cosmology: $\Omega_{\Lambda}=0.6911$, $\Omega_{m}=0.3089$, $\Omega_{b}=0.0486$, $\sigma_{8}=0.8159$, $n_{s}=0.9667$, and $h=0.6774$ (Planck Collaboration et al., 2016).
For this work, we use the largest volume, TNG300, which is 205 $h^{-1}$ Mpc on each side. TNG300 has three different realizations, with TNG300-1 having $2500^{3}$ DM particles, TNG300-2 having $1250^{3}$ DM particles, and TNG300-3 $625^{3}$ DM particles. We use the $z=0$ snapshot of TNG300-3, but there are 127 snapshots going back to $z=127$ (with the same redshifts as the original Illustris simulation). This simulation has a mass resolution of $M_{\rm DM}=4.5\times 10^{9}~{}M_{\odot}$. The proof-of-concept models described in this paper are trained on either the spatial distribution of DM particle mass density or that of the DM subhalos. Since we are only interested in the distribution of matter on the largest scales, we determined that the DM-only versions of the TNG300 volumes were sufficient for our purposes. In the future, as we continue to enhance our void detectors, we plan to incorporate additional information from simulations, such as galaxy luminosity and color, to improve predictive performance.
To train on data sets that emulate the sparseness of galaxy redshift surveys, we use DM halo samples with number densities comparable to the Sloan Digital Sky Survey (SDSS; Abazajian et al., 2009a) main galaxy sample or to the DESI Bright Galaxy Survey (e.g., Hahn et al., 2023b). We use the subhalo catalog that was generated using the SUBFIND algorithm (Springel et al., 2001), a computational tool to identify and analyze substructures in DM halos. These subhalos are smaller gravitationally-bound clumps of matter within each halo.
3 Tidal Tensor Definition
Figure 1: Left: TNG300-3-Dark mass density contrast defined as $\delta(\vec{x})=\rho(\vec{x})/{\bar{\rho}}-1$, where $\bar{\rho}$ is the average matter density in the volume. Center: corresponding gravitational potential $\Phi$. Note that both $\delta$ and $\Phi$ have been scaled and shifted such that their minimum value is zero and their standard deviation is one, and their logarithm plotted for ease of interpretability. Right: multi-label tidal tensor classified mask. This mask is our “truth table” for training. These plots are one slice thick, with $N_{\rm mesh}=512$, so they are 0.4 $h^{-1}$ Mpc thick.
The best definition of cosmic voids and the optimal method for finding them may depend on the astrophysical or cosmological question of interest. For example, if the subject of interest is the difference in evolution between void and non-void galaxies, how well the definition delineates between void and wall regions is of paramount importance (Douglass et al., 2023; Veyrat et al., 2023; Zaidouni et al., 2024). On the other hand, if one wants to investigate cosmological effects such as the impact of voids on the CMB due to lensing or through the ISW effect, then obtaining accurate locations of void centers with respect to the potential is more important (Cai et al., 2017). Here we adopt a specific well-motivated physical definition. In future work, models could be trained using the same deep-learning network architecture applied to different physical definitions.
For training our proof-of-concept deep learning void detector, we employ a physical definition of voids that is based on simple physics: the gravitational potential that drives the formation of large-scale structure. Thus, we define structural morphology with the so-called “T-web” method, which was first used to segment large-scale structures by Hahn et al. (2007) (see also Forero-Romero et al., 2009). This method is based on the tidal tensor, the Hessian of the gravitational potential:
Hahn et al. (2007) then found the eigenvalues of the tidal tensor and counted the number of positive eigenvalues to classify the environment of each voxel. Motivated by Zel’dovich theory (Zel’dovich, 1970), they defined voids as regions where $T_{ij}$ has no positive eigenvalues (unstable orbits), walls as regions with one positive eigenvalue (one-dimensional stable manifold), filaments having two positive eigenvalues (two-dimensional stable manifold), and finally halos as regions with all three positive eigenvalues (attractive fixed points). The tidal tensor is a good definition for our purposes because it captures the local gravitational dynamics in a region in an intuitive and relatively simple way. Furthermore, the T-web method uses only two free parameters: the smoothing scale, $\sigma$, and the eigenvalue threshold, $\lambda_{\mathrm{th}}$, used for distinguishing between the classes.
Related methods of classifying structural morphology include use of the Hessian of the density field (Aragón-Calvo et al., 2007), use of the Hessian of the velocity shear tensor (the “V-web method”; Hoffman et al., 2012), and ORIGAMI (Falck et al., 2012), which counts stream crossings that have already occurred. Libeskind et al. (2018) review and compare methods of identifying structural elements of the cosmic web. Among their findings is that agreement between methods is highest for void regions (as opposed to denser structures, where some methods substantially disagree). The deep-learning architecture we use, as described in Section 4, could be applied to other methods of constructing the truth table for training, so future versions of DeepVoid could be trained to leverage different definitions of the morphology for different cosmological purposes.
To perform a classification using the tidal tensor, we must calculate the gravitational potential $\Phi$. First, we compute the matter density contrast $\delta(\vec{r})$ by using the cloud-in-cell (CIC) interpolation scheme (Hockney & Eastwood, 1988) to allocate each DM particle’s mass appropriately in a grid with $N_{\rm mesh}^{3}$ voxels, where $N_{\rm mesh}$ is the number of voxels on a side. Then, by using fast Fourier transforms (FFTs) and a Green’s function to solve Poisson’s equation,
$\hat{\Phi}(\vec{k})=-4\pi G\frac{\hat{\rho}(\vec{k})}{|\vec{k}|^{2}},$ (2)
we obtain the gravitational potential $\hat{\Phi}(\vec{k})$ in Fourier space, where $\vec{k}$ is the wavevector and hats denote values in Fourier space. See Appendix A for more details on calculating the tidal tensor. We can then inverse-FFT this to obtain the potential in the spatial domain, $\Phi(\vec{x})$.
We smooth the gravitational potential, $\Phi$, with a Gaussian kernel of width $\sigma_{\rm eff}$, where $\sigma_{\rm eff}^{2}=\sigma^{2}+\sigma_{G}^{2}$. This allows us to take into account the intrinsic smoothing that comes with defining a grid of a certain size, where $\sigma_{G}=L_{\rm box}/N_{\rm mesh}$ is the grid size. All potentials in this work are smoothed with $\sigma_{\rm eff}=1$ $h^{-1}$ Mpc for consistency. For the TNG300 simulation, we choose a grid with $N_{\rm mesh}=512$, so each voxel is 0.4 $h^{-1}$ Mpc on a side.
@@ 3 Tidal Tensor Definition
Figure 2: A comparison of different eigenvalue thresholds used to create a tidal tensor-based map of structural morphology in TNG300-3-Dark. Top left: $\lambda_{\mathrm{th}}=0$; Top right: $\lambda_{\mathrm{th}}=0.3$; Bottom left: $\lambda_{\mathrm{th}}=0.65$ (our default); Bottom right: $\lambda_{\mathrm{th}}=1$. The morphology of the cosmic web as defined by $\lambda_{\mathrm{th}}$$=0$ consists of tiny isolated voids surrounded by huge amounts of wall voxels. We see a more realistic depiction of the Universe’s large-scale structure at slightly positive values of $\lambda_{\mathrm{th}}$.
@@
One notable change to the method of Hahn et al. (2007) is our use of a non-zero eigenvalue threshold $\lambda_{\mathrm{th}}$, as recommended in previous studies using similar techniques (Forero-Romero et al., 2009; Alonso et al., 2015). If we apply the exact procedure of Hahn et al. (2007) to the TNG300 sample, we do not obtain a classification of the large-scale structure that matches what we observe in galaxy surveys or expect in simulations. If, as they did, we choose $\lambda_{\mathrm{th}}=0$ to differentiate between compressing and lengthening tidal forces, we find that only 16% of the total TNG300 volume is classified as void.
Choosing $\lambda_{\mathrm{th}}$$=0$ means that if an eigenvalue is even infinitesimally positive, the classification assumes that the locality around that grid point collapses along that eigenvector. This collapse happens over a dynamical timescale. Thus, for an infinitesimally positive value of the eigenvalue, collapse may either never occur (if the value is positive due to noise or transient behavior) or not until the remote future (possibly larger than the Hubble time). As Forero-Romero et al. (2009) notes, the dimensionality of $T_{ij}$ is [time]-2, so its eigenvalues can be correlated with the free-fall collapse time:
@@
where $\beta(\lambda_{i})$ is a corrective factor that accounts for deviations from perfect isotropy. If we use TNG’s Planck 2015 cosmological parameters and assume isotropy, $\beta(\lambda_{i})=1$, then this order-of-magnitude calculation suggests a threshold for voids of $\lambda_{\mathrm{th}}=2.58$. Alonso et al. (2015) argue that, because the density contrast is the sum of the eigenvalues, and non-linear collapse occurs around a density contrast of unity, this eigenvalue threshold should be of order unity. As other authors who have used the tidal tensor method discuss (Forero-Romero et al., 2009; Alonso et al., 2015; Bonnaire et al., 2022), the choice of value for this parameter depends on the properties of the cosmic web that are of interest. For example, if we were only concerned with finding the most underdense core regions of voids, then a threshold of 0 would be useful.

### v2

@@ Abstract
We present DeepVoid, an application of deep learning trained on a physical definition of cosmic voids to detect voids in density fields and galaxy distributions. By semantically segmenting the IllustrisTNG simulation volume using the tidal tensor, we train a deep convolutional neural network to classify local structure using a U-Net architecture for training and prediction. The model achieves a void F1 score of 0.96 and a Matthews correlation coefficient over all structural classes of 0.81 for dark matter particles in IllustrisTNG with interparticle spacing of $\lambda=0.33$ $h^{-1}$ Mpc. We then apply the machine learning technique of curriculum learning to enable the model to classify structure in data with significantly larger intertracer separation. At the highest tracer separation tested, $\lambda=10$ $h^{-1}$ Mpc, the model achieves a void F1 score of 0.89 and a Matthews correlation coefficient of 0.6 on IllustrisTNG subhalos.
I Introduction
The large-scale structure of the Universe, as observed in galaxy redshift surveys, resembles a weblike structure (Bond et al., 1996). This structure is, by volume, chiefly comprised of cosmic voids, underdense regions that span tens of megaparsecs and are bounded by sheets and filaments of galaxies that flow into large galaxy clusters. Depending on the definition, voids fill approximately $\sim$70% of the volume of the Universe (Pan et al., 2012). Voids were found in some of the earliest galaxy and cluster redshift surveys (e.g., Gregory and Thompson, 1978; Jõeveer et al., 1978) and were shown to be ubiquitous features of the large-scale structure by dense three-dimensional galaxy redshift surveys (de Lapparent et al., 1986; Geller and Huchra, 1989; Shectman et al., 1996).
Voids represent an interesting opportunity for cosmology and physics (Cai and Neyrinck, 2025). Voids grow from initial troughs in the primordial field of density fluctuations and continue to become more vacuous as the large-scale structure in the Universe evolves. They are ideal laboratories for studying modifications to general relativity (Perico et al., 2019) by observing, for example, the weak lensing signature of voids (Baker et al., 2018) or the gravitational redshift of cosmic microwave background (CMB) photons by large voids, known as the integrated Sachs-Wolfe (ISW) effect (Cai et al., 2014).
Investigations of dark energy (DE) can also greatly benefit from studying voids. Because overdense structures such as galaxy clusters formed long before dark energy could strongly influence the evolution of large-scale structure, voids are the primary component of the cosmic web influenced by DE. Voids act as super-Hubble bubbles that expand faster than the Hubble flow (Goldberg and Vogeley, 2004). To leading order, the accelerated expansion of voids is the primary signature of DE; the overdense structures accelerate away from each other because the voids are expanding between them (Sheth and van de Weygaert, 2004).
Several specific cosmological tests using cosmic voids are of particular interest and motivate how our void finder is designed and implemented. The shapes of voids provide an avenue to test cosmological parameters through the Alcock-Paczyński (AP) effect (e.g., Alcock and Paczynski, 1979; Lavaux and Wandelt, 2012; Hamaus et al., 2020). The size spectrum of voids is sensitive to both the growth of structure and the DE equation of state (e.g., Nadathur, 2016; Verza et al., 2019; Contarini et al., 2022, 2023). Observation of the ISW effect from voids can further constrain cosmological models (e.g., Granett et al., 2008; Kovács, 2018).
One of the chief difficulties in finding and using voids for cosmology and astrophysics is the definition of a void. There is no universally agreed-upon definition, but studies generally rely on finding underdensities in the galaxy distribution, with methods mostly varying in the shape of the density estimation kernel. Popular void-finding methods for galaxy surveys utilize techniques that include the detection of voids as spherical regions or unions of multiple spheres (e.g., Kauffmann and Fairall, 1991; El-Ad and Piran, 1997; Hoyle and Vogeley, 2002; Banerjee and Dalal, 2016; Zhao et al., 2016; Paz et al., 2023), and watershed segmentation of the galaxy density field (e.g., Platen et al., 2007; Neyrinck, 2008; Sutter et al., 2015; Nadathur et al., 2019). Void catalogs constructed from identical galaxy samples can differ drastically based on how a void is defined in that study (Colberg et al., 2008). These different definitions make it complicated to compare analyses based on different void catalogs. Many of these methods also rely to varying degrees on fine-tuned parameters and different a priori notions of what constitutes a void and how they should appear.
In this work, we demonstrate a proof of concept of an alternative approach: define voids in terms of the physics that creates and shapes them, and train a deep neural network to recognize the spatial signature of that physical process. Here we describe DeepVoid, a void detector based on physical criteria to identify voids using convolutional neural networks (CNNs). By employing such criteria, we can train CNNs to detect voids that are useful for cosmological tests of interest. Previous investigations of using deep learning for morphological classification include Aragon-Calvo (2019) and Inoue et al. (2022). Unlike previous work, DeepVoid is targeted to the specific task of detecting voids in galaxy redshift surveys (although, as we will show, it can identify all the components of the cosmic web), a task which grows more challenging as these surveys probe deeper into the Universe.
The eventual goal of this approach is to fully leverage the cosmological information available in current and upcoming galaxy surveys and instruments, such as the Dark Energy Spectroscopic Instrument (DESI; DESI Collaboration et al., 2016), the Vera C. Rubin Legacy Survey of Space and Time (LSST; LSST Science Collaboration et al., 2009), Euclid (Euclid Collaboration et al., 2024), the Nancy Grace Roman Space Telescope (Schlieder et al., 2024), and the Subaru/Hyper Suprime Camera (HSC; Miyazaki et al., 2018). These surveys and others will provide maps of large-scale structure to unprecedented depths and with remarkable precision. They will probe ranges of cosmological redshift over which the evolution of voids is apparent and over which details of DE properties, possible modifications to general relativity, and neutrino properties might be probed (Massara et al., 2015; Pisani et al., 2015; Perico et al., 2019; Avsajanishvili et al., 2024).
Important challenges include that the galaxy catalogs from these new surveys may be quite sparse compared to more local galaxy surveys and that large-scale structure is probed by many different types of galaxy tracers with different biases (Euclid Collaboration et al., 2022; Wang et al., 2022; Hahn et al., 2023; Zhou et al., 2023; Raichoor et al., 2023; Chaussidon et al., 2023). The effect of “tracer bias” on voids has been shown to affect the void size distribution (Nadathur and Hotchkiss, 2015; Contarini et al., 2019), impact the tracer profile vs. matter profile in voids (Pollina et al., 2017), and alter the velocity profile in voids, which is important for using void-galaxy correlations to study redshift-space distortions (RSDs; Massara et al., 2022). Current methods of estimating the void spectrum involve corrections for that bias applied after making a void catalog, based on analysis of mock galaxy catalogs (Contarini et al., 2022). We will minimize the need for such postdetection void corrections by using mock catalogs to train variants of DeepVoid for specific galaxy target classes. In this proof-of-concept paper, we demonstrate an approach to tackling the sparse sampling problem. In a follow-up paper, we will examine how the deep learning approach addresses the variety of galaxy tracers.
In this work, we present DeepVoid, a novel multiclass method of segmenting the morphology of structures in cosmological simulations using CNNs trained to match the classes assigned from properties of the gravitational potential. This paper is organized as follows. In Section II, we describe the simulations from which we define our truth table. The physical definition of a void used to generate the truth table is discussed in Section  III. Aspects of deep learning and the U-Net architecture are detailed in Section IV. Model training and performance are presented in Sections V and VI, respectively. Training and performance of models that predict on sparse samples of halos are described in Section VII. Conclusions and thoughts about future work are shared in Section VIII. Additional details on multiclass performance metrics, our use of batch normalization, and different loss functions can be found in the Appendices.
II Simulation Data
The work presented here is based on The Next Generation Illustris Simulation (TNG) (Springel et al., 2018; Pillepich et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Marinacci et al., 2018). TNG is a gravo-magnetohydrodynamical simulation run using the moving-mesh code AREPO (Springel, 2010; Vogelsberger et al., 2014). The TNG suite is made up of three volumes of different sizes: 35, 75, and 205 $h^{-1}$ Mpc on a side. All TNG runs were computed with the Planck 2015 cosmology: $\Omega_{\Lambda}=0.6911$, $\Omega_{m}=0.3089$, $\Omega_{b}=0.0486$, $\sigma_{8}=0.8159$, $n_{s}=0.9667$, and $h=0.6774$ (Planck Collaboration et al., 2016).
For this work, we use the largest volume, TNG300, which is 205 $h^{-1}$ Mpc on each side. TNG300 has three different realizations, with TNG300-1 having $2500^{3}$ DM particles, TNG300-2 having $1250^{3}$ DM particles, and TNG300-3 having $625^{3}$ DM particles. We use the $z=0$ snapshot of TNG300-3, but there are 127 snapshots going back to $z=127$ (with the same redshifts as the original Illustris simulation). The proof-of-concept models described in this paper are trained on either the spatial distribution of DM particle mass density or that of the DM subhalos. Since we are only interested in the distribution of matter on the largest scales, we determined that the DM-only versions of the TNG300 volumes were sufficient for our purposes. TNG300-3-DARK has a mass resolution of $M_{\rm DM}=3\times 10^{9}h^{-1}~M_{\odot}$. In the future, as we continue to enhance our void detectors, we plan to incorporate additional information from simulations, such as galaxy luminosity and color, to improve predictive performance.
To train on data sets that emulate the sparseness of galaxy redshift surveys, we use DM halo samples with number densities comparable to the Sloan Digital Sky Survey (SDSS; Abazajian et al., 2009) main galaxy sample or to the DESI Bright Galaxy Survey (e.g., Hahn et al., 2023). We use the subhalo catalog that was generated using the SUBFIND algorithm (Springel et al., 2001), a computational tool to identify and analyze substructures in DM halos. These subhalos are smaller gravitationally-bound clumps of matter within each halo.
III Tidal Tensor Definition
Figure 1: Left: TNG300-3-Dark mass density contrast defined as $\delta(\vec{x})=\rho(\vec{x})/{\bar{\rho}}-1$, where $\bar{\rho}$ is the average matter density in the volume. Center: corresponding gravitational potential $\Phi$. Note that both $\delta$ and $\Phi$ have been scaled and shifted such that their minimum value is zero and their standard deviation is one, and their logarithm plotted for ease of interpretability. Right: multilabel tidal tensor classified mask. This mask is our “truth table” for training. These plots are one slice thick, with $N_{\rm mesh}=512$, so they are 0.4 $h^{-1}$ Mpc thick.
The best definition of cosmic voids and the optimal method for finding them may depend on the astrophysical or cosmological question of interest. For example, if the subject of interest is the difference in evolution between void and nonvoid galaxies, how well the definition delineates between void and wall regions is of paramount importance (Douglass et al., 2023; Veyrat et al., 2023; Zaidouni et al., 2024). On the other hand, if one wants to investigate cosmological effects such as the impact of voids on the CMB due to lensing or through the ISW effect, then obtaining accurate locations of void centers with respect to the potential is more important (Cai et al., 2017).
For training our proof-of-concept deep learning void detector, we employ a definition of voids that is based on simple physics: the gravitational potential that drives the formation of large-scale structure. Thus, we define structural morphology with the so-called “T-web” method, which was first used to segment large-scale structures by Hahn et al. (2007) (see also Forero-Romero et al., 2009). This method is based on the tidal tensor, the Hessian of the gravitational potential:
Hahn et al. (2007) then found the eigenvalues of the tidal tensor and counted the number of positive eigenvalues to classify the environment of each voxel. Motivated by Zel’dovich’s theory (Zel’dovich, 1970), they defined voids as regions where $T_{ij}$ has no positive eigenvalues (unstable orbits), walls as regions with one positive eigenvalue (one-dimensional stable manifold), filaments having two positive eigenvalues (two-dimensional stable manifold), and finally halos as regions with all three positive eigenvalues (attractive fixed points). The tidal tensor is a good definition for our purposes because it captures the local gravitational dynamics in a region in an intuitive and relatively simple way. Furthermore, the T-web method uses only two free parameters: the smoothing scale, $\sigma$, and the eigenvalue threshold, $\lambda_{\mathrm{th}}$, used for distinguishing between the classes.
For the purpose of identifying voids, the tidal tensor definition is particularly useful because it describes the spatial pattern of void regions around peaks of the potential. By using the tidal tensor as the truth table upon which the network is trained in the DeepVoid model described below, voids are identified not simply as regions of low density, but also based on the spatial pattern of matter that is predicted to arise from gravity.
The tidal tensor definition of structural segmentation has been widely applied to simulations of large-scale structure, in which one has access to the exact matter density field, and so one can compute the gravitational potential and its derivatives using fast Fourier methods. This definition is difficult to apply to observational data because the gravitational potential is not directly observable and depends on long-wavelength modes of the matter density field. Estimation of the potential and its derivatives from galaxy survey data is further limited by the survey boundaries, sparseness of tracers of the matter density, biasing of tracers, and RSDs. Wang et al. (2012) estimate the tidal tensor within the SDSS main galaxy sample region and find acceptable results only very far from the survey boundaries. Eardley et al. (2015) examine how estimation of the tidal tensor is influenced by the survey geometry, galaxy sampling, and redshift space distortions in application to the GAMA survey.
Related methods of classifying structural morphology include the use of the Hessian of the density field (Aragón-Calvo et al., 2007), the use of the Hessian of the velocity shear tensor (the “V-web method”; Hoffman et al., 2012), and ORIGAMI (Falck et al., 2012), which counts stream crossings that have already occurred. Libeskind et al. (2018) review and compare methods of identifying structural elements of the cosmic web. Among their findings is that agreement between methods is highest for void regions (as opposed to denser structures, where some methods substantially disagree). The deep learning architecture we use, as described in Section IV, could be applied to other methods of constructing the truth table for training, so future versions of DeepVoid could be trained to leverage different definitions of the morphology for different cosmological purposes.
To perform a classification using the tidal tensor, we must calculate the gravitational potential $\Phi$. First, we compute the matter density contrast $\delta(\vec{r})$ by using the cloud-in-cell (CIC) interpolation scheme (Hockney and Eastwood, 1988) to allocate each DM particle’s mass appropriately in a grid with $N_{\rm mesh}^{3}$ voxels, where $N_{\rm mesh}$ is the number of voxels on a side. Then, by using fast Fourier transforms (FFTs) and a Green’s function to solve Poisson’s equation,
$\hat{\Phi}(\vec{k})=-4\pi G\frac{\hat{\rho}(\vec{k})}{|\vec{k}|^{2}},$ (2)
we obtain the gravitational potential $\hat{\Phi}(\vec{k})$ in Fourier space, where $\vec{k}$ is the wavevector and hats denote values in Fourier space. We can then inverse-FFT this to obtain the potential in the spatial domain, $\Phi(\vec{x})$. In Appendix A we provide more mathematical details of the Fourier method for computing the potential.
@@ III Tidal Tensor Definition
Figure 2: A comparison of different eigenvalue thresholds used to create a tidal-tensor-based map of structural morphology in TNG300-3-Dark. Top left: $\lambda_{\mathrm{th}}=0$; Top right: $\lambda_{\mathrm{th}}=0.3$; Bottom left: $\lambda_{\mathrm{th}}=0.65$ (our default); Bottom right: $\lambda_{\mathrm{th}}=1$. The morphology of the cosmic web as defined by $\lambda_{\mathrm{th}}$=0=0 consists of tiny isolated voids surrounded by huge amounts of wall voxels. We see a more realistic depiction of the Universe’s large-scale structure at slightly positive values of $\lambda_{\mathrm{th}}$.
@@
One notable change to the method of Hahn et al. (2007) is our use of a nonzero eigenvalue threshold $\lambda_{\mathrm{th}}$, as recommended in previous studies using similar techniques (Forero-Romero et al., 2009; Alonso et al., 2015). If we apply the exact procedure of Hahn et al. (2007) to the TNG300 sample, we do not obtain a classification of the large-scale structure that matches what we observe in galaxy surveys or expect in simulations. If, as they did, we choose $\lambda_{\mathrm{th}}=0$ to differentiate between compressing and lengthening tidal forces, we find that only 16% of the total TNG300 volume is classified as void.
Choosing $\lambda_{\mathrm{th}}$=0=0 means that if an eigenvalue is even infinitesimally positive, the classification assumes that the locality around that grid point collapses along that eigenvector. This collapse happens over a dynamical timescale. Thus, for an infinitesimally positive value of the eigenvalue, collapse may either never occur (if the value is positive due to noise or transient behavior) or not until the remote future (possibly larger than the Hubble time). As Forero-Romero et al. (2009) note, the dimensionality of $T_{ij}$ is [time]-2, so its eigenvalues can be correlated with the freefall collapse time:
@@
where $\beta(\lambda_{i})$ is a corrective factor that accounts for deviations from perfect isotropy. If we use TNG’s Planck 2015 cosmological parameters and assume isotropy, $\beta(\lambda_{i})=1$, then this order-of-magnitude calculation suggests a threshold for voids of $\lambda_{\mathrm{th}}=2.58$. Alonso et al. (2015) argue that, because the density contrast is the sum of the eigenvalues, and nonlinear collapse occurs around a density contrast of unity, this eigenvalue threshold should be of order unity. As other authors who have used the tidal tensor method discuss (Forero-Romero et al., 2009; Alonso et al., 2015; Bonnaire et al., 2022), the choice of value for this parameter depends on the properties of the cosmic web that are of interest. For example, if we were only concerned with finding the most underdense core regions of voids, then a threshold of zero would be useful.